### v5

@@ 3 Empirical Approach
$u_{i,t}$ $=\left(\boldsymbol{\beta}_{i}\boldsymbol{\theta}_{i}^{x}+\boldsymbol{\theta}_{i}\right)\boldsymbol{\omega}_{t}+\left[\boldsymbol{\beta}\left(\boldsymbol{\gamma}_{i}^{x}+\boldsymbol{\theta}_{i}^{x}\boldsymbol{\xi}\right)\right]+\boldsymbol{\gamma}_{i}+\boldsymbol{\theta}_{i}\boldsymbol{\xi}+\epsilon_{i,t}.\boldsymbol{F}_{t}\boldsymbol{\phi}_{i}\boldsymbol{\omega}_{i,t}+\boldsymbol{\epsilon}_{i,t}.$
@@
Yet, our aim is the identification of the interaction of the observed common factors and the country-specific variation $\boldsymbol{\omega}_{i,t}$. To identify $\hat{u}_{i,t}$ correctly, we must ensure that no common factors or cross-section dependence is left in $\epsilon_{i,t}$.
@@
In the second step, we therefore extract the common components of the observed common factors, of the unobserved ones, and of the remaining residual variation from $\boldsymbol{\hat{u}}_{i,t}$ by using principal components. The number of components is determined by the criteria of Ahn and Horenstein (2013). In the third step, we then replace the observed and unobserved common factors with the principal components from the previous step and estimate the following model:
@@
where $\boldsymbol{PC}_{t}$ are the first $p_{pc}$ principal components of the observed and approximated common factors and of the estimated residuals. [^15] Our model is related to the regularized CCE estimator in Juodis (2022). The idea is to reduce the dimension of the cross-section averages to the dimension of the estimated number of common factors.
@@ 4.1 Benchmark Results
Real Interest Rate -0.038 -0.013 -0.019 -0.182 -0.311 -0.070 -0.070
@@
REER 0.098 0.106 0.066 0.700 0.664 0.424 0.424
(0.025)*** (0.025)*** (0.026)** (0.284)** (0.256)*** (0.187)** (0.187)**
Capital-to-Asset -0.217 -0.216 -0.264 -0.131 -0.290 0.815 0.815
@@
This table displays the results for the years 2004-2022 with the non-core ratio as dependent variable. Standard errors in parenthesis. Significance levels: ${}^{*}10\%,^{**}5\%,^{***}1\%$. TWFE and TWFE-I include country and time fixed effects. TWFE-I additionally includes interactions between the fixed exchange rate dummy and time fixed effects. All explanatory variables except Capital Regulation and Liquidity Regulation are added as cross-section averages for MG-CCE. Principal Components (PC) are extracted from composite $\hat{u}_{i,t}$ in column MG-CCE. CD is the value of the CD test (Pesaran, 2015) with the null hypothesis of weak cross-section dependence. $\hat{M}_{X}$ is the estimate of the number of common factors using the Growth Criterion from Ahn and Horenstein (2013), with $\hat{u}$ being the composite of observed and unobserved common factors and $\hat{\epsilon}$ the residual. All variables are defined as in Table 9.
@@
Additional Results corresponding to columns (4), (5), (6) and (7) of Table 2. $F:PC=0$ tests if the coefficients of the principal components are joint significantly different from zero. $\Delta(all)$ test the null of slope homogeneity of all coefficients, $\Delta(PC)$ test the null of slope homogeneity of the coefficients of the principal components under the assumption that all other coefficients are heterogeneous. $PCF_{1}>PCF_{0}$ test if the sum of the coefficients of the PCs for $fixed=1$ are larger than those for $fixed=0$. Share PC is the share of the explained variance of the PC. All variables are defined as in Table 9.
@@
The first column of Table 2 presents a standard two-way fixed effects specification with homogeneous coefficients, country and year fixed effects and no accounts of any observed or unobserved common factors. [^17] We use Stata 18 and the community contributed packages reghdfe (Correia, 2023) for the TWFE and FE estimations and xtdcce2 and xtcd2 (Ditzen, 2018, 2021) for the MG-CCE regressions and the CD tests. The number of common factors is estimated using xtnumfac (Ditzen and Reese, 2023). The CD test rejects the null of weak cross-section dependence, i.e., it suggests that there are global factors that are not accounted for and/or those global factors are affecting countries in a distinct way. [^18] Juodis and Reese (2021) show that the CD test statistic diverges under certain conditions and propose a corrected test. Our results are qualitatively similar using the their test. This is supported by two estimated common factors in the residuals ($\hat{M}_{\hat{\epsilon}}$). Moving rightwards along the table, we attempt to account for the factor structure in two ways. In column (2), we add the observed common factors and in column (3), we interact the time fixed effects with the fixed exchange rate indicator. Two of the observed common factors are now significant, but even in this case, strong cross-section dependence remains in the residuals, indicating a potential bias of the regressions results.
@@
Moving to columns (4) and (5), we change our specification in two aspects. First, we allow for country-specific coefficients on both the country-specific variables and observed global factors using the mean group estimator (MG). The test for slope heterogeneity, presented in the columns labelled (4) and (5) in Table 3 confirms the necessity to allow for country-specific slopes. [^19] The tests for slope heterogeneity were performed with the xthst package (Bersvendsen and Ditzen, 2021). Turning to the results in column (4) of Table 2, while no factors remain in the residuals, the CD test still rejects the null of weak cross-section dependence. Further, none of the observed common factors are significant. While this specification potentially accounts for the factor structure, the results remain questionable due to the strong cross-section dependence. To overcome this weakness, in column (5), we then employ the CCE estimator as per equation (3)—stage (1) of our three step procedure—and add, in addition to the observed common factors, cross-section averages to fully account for the unobserved factor structure. [^20] We add the cross-section averages of the dependent variable, real exchange rate, GDP growth, the real interest rate, capital-to-asset ratio and money over GDP. We do not add the cross-section averages of the macroprudential variables as they are part of the observed common factor ’World Macroprudential’. The six cross-section averages ensure that the rank condition holds, i.e., the number of (un)observed factors is smaller or equal to the number of cross-section averages, see Karabiyik et al. (2017), DeVos et al. (2024). The CD-test statistic falls into a non-rejection area (CD=0.798) and the estimate of the number of common factors in the residual is zero so that we can now calculate the composite $\hat{u}_{i,t}$, which encapsulates the entire factor structure. As we found three common factors in the non-core to core ratio, in the second step, we extract the first three principal components (PCs) from the composite $\hat{u}_{i,t}$. The PCs account for roughly 80% of the explained variance, see the lower panel in Table 3.212121We note that the estimate for the number of factors in the composite, $\hat{M}_{\hat{u}}$, is 2, the same as the number of factors in the residuals of the (TW)FE regressions. No factors remain in the residual, so we conclude that the composite captures the entire factor structure. Columns (6) and (7) present the results of the 3rd stage. Introducing the PCs successfully mitigates cross-sectional dependence, as indicated by the CD-statistic of -0.089 (p=0.929). As explained in the previous section, because the estimated PCs span both observed and unobserved global factors, their inclusion in the MG regression makes the observed common factors redundant (and display high multicollinearity with some of them as discussed below), so the latter are omitted. Hence we conclude that these two specifications fully account for the factor structure. In column (6), all three PCs are significant and positive, implying an increase in the non-core ratio following a shock propagated by the PCs. Column (7) then tests whether global factors (captured by the three PCs) affect the non-core ratio of countries on a fixed exchange rate more than those on a floating exchange rate. The loadings for the first PC is higher in countries subject to fixed exchange rates (0.993 vs. 0.304) and significant. PC2 is positive and significant for fixers, while PC3 is significant for floaters. In general the loadings for fixers are larger than those for floaters, as confirmed by the test statistic of 5.18 on the differences between the sum of the PC loadings between fixed and flexible exchange rate regimes, shown in the column labelled (7) in Table 3. Therefore, the upshot is that there is cause for the trilemma hypothesis applied to banks’ non-core funding ratios.
@@
Across all specifications, it is mainly three domestic variables that affect banks’ non-core ratios in a statistically significant manner. First, the lagged dependent variable is positive and significant at the 1% level throughout. Second, banking sectors in countries with a more appreciated real exchange rate have higher non-core ratios. This result might reflect that real exchange rates appreciate in countries with a higher domestic interest rate, which in turn incentivizes banks to borrow abroad. Finally, higher money-to-GDP ratios also tend to increase banks’ non-core ratios, consistent with the evidence on drivers of credit booms and the role of cross-border borrowing found in the studies referred to in previous sections.
@@ 4.2 Robustness Checks
In this sub-section, we show that our main results corresponding to column (7) of Table 2 are robust along several dimensions. We first drop the United States from our sample, as most of the observed common factors included in the regressions are based on (or at least dominated by) the US. Our benchmark results are hardly affected and we still find PC1 and PC3 to affect floaters, and PC1 and PC2 to affect fixers to a statistically significant extent, as column (1) of Table 4 shows.
In column (2), we use our benchmark sample including Iceland, which we dropped for the benchmark analysis because it is an extreme outlier for several of the variables included in the analysis. As becomes apparent, most results are similar to our baseline results, but the coefficients are estimated more imprecisely, presumably due to the additional uncertainty introduced by the inclusion of Iceland. In addition, the specification is left with some cross-sectional dependence.
@@
Finally, in column (5), we use an alternative outcome variable—banks’ loan-to-deposit ratio. Previous work suggested it to be closely correlated with the non-core to core ratio, as we also confirm for our sample in Section 2. For this outcome variable, we find PC1 and PC2 to be significant drivers in countries with a floating exchange rate and PC1 and PC3 to be statistically significant for countries on a fixed exchange rate. One difference to the previous results that we obtained for banks’ non-core ratios is that a flexible exchange rate does not provide any insulation to PC1. This result, however, is driven by the crises episodes included in column (5). In fact, when in unreported specifications we drop the Covid episode and/or the global financial crisis from the sample, we find that the loadings to PC1 are substantially larger for fixers than for floaters, as in our benchmark analysis.
@@
Real Interest Rate -0.097 -0.095 -0.087 0.066 -0.088
@@
This table displays the robustness checks. The non-core ratio is the dependent variable in Columns (1)-(4). In Column (5), LTD, the loan-to-deposit ratio, is the dependent variable. Standard errors in parenthesis. Significance levels: ${}^{*}10\%,^{**}5\%,^{***}1\%$. Column Excl. US excludes the United States. Column Inc. Iceland includes Iceland. Column No Covid spans the time period 2004:Q1-2019:Q4 and No Crisis 2010:Q1-2019:Q4. The 1st stage for each column is not displayed. The PCs are obtained from this initial stage. $\hat{M}_{X}$ is the estimate of the number of common factors using the Growth Criterion from Ahn and Horenstein (2013), with $\hat{u}$ being the composite of observed and unobserved common factors and $\hat{\epsilon}$ the residual. All variables are defined as in Table 9. For further details see notes of Table 2.
@@ 4.3 Does Macroprudential Policy Insulate Countries from the GFC?
(1) (2) (3)
All Fixed = 0 Fixed = 1
Lag Dep. Var. 0.659 0.701 0.623
(0.031)*** (0.045)*** (0.041)***
GDP Growth -0.015 -0.002 -0.026
(0.091) (0.042) (0.169)
Real Interest Rate -0.070 0.123 -0.239
(0.072) (0.069)* (0.105)**
REER 0.424 0.079 0.726
(0.187)** (0.036)** (0.335)**
Capital-to-Asset 0.815 0.554 1.044
(0.486)* (0.615) (0.750)
Capital Reg. -4.003 -0.282 -7.259
(2.899) (0.824) (5.333)
Liquidity Reg. -0.295 -0.116 -0.452
(2.976) (0.961) (5.604)
Money / GDP 0.175 0.081 0.257
(0.064)*** (0.053) (0.109)**
Principal Components (Macro. Tight. = 0)
PC1 0.413 0.138 0.654
(0.469) (0.271) (0.856)
PC2 0.786 0.020 1.457
(0.307)** (0.096) (0.519)***
PC3 0.937 0.227 1.558
(0.550)* (0.177) (1.009)
Principal Components (Macro. Tight. = 1)
PC1 0.884 0.513 1.208
(0.305)*** (0.220)** (0.535)**
PC2 -0.083 -0.195 0.015
(0.244) (0.326) (0.366)
PC3 0.260 0.575 -0.015
(0.291) (0.402) (0.415)
Obs 2160 1008 1152
Countries 30 14 16
Time Periods 72 72 72
CD -0.089 2.100 0.920
p 0.929 0.036 0.357
$\hat{M}_{\hat{\epsilon}}$ 0 0 0
$R^{2}$ 0.042 0.057 0.041
$R^{2}$ (MG) 0.947 0.929 0.949
$R^{2}$ (within) 0.958 0.943 0.959
1st Stage
CD 0.798 0.798 0.798
p 0.425 0.425 0.425
$\hat{M}_{\hat{u}}$ 2 2 2
$\hat{M}_{\hat{\epsilon}}$ 0 0 0
@@
Principal Components obtained from Baseline Regression in Table 2. The regressions in this table identify the loadings to the PCs separately for countries with loose vs. tight macroprudential policies. Column Fixed = 0 restricts the sample to countries with floating exchange rates, Fixed = 1 to countries with fixed exchange rates. See the additional notes of Table 2. All variables are defined as in Table 9. For further details see notes of Tables 2 and 4.
Table 5: The Relevance of Macroprudential Policies
Previous regressions have shown that a floating exchange rate helps insulate countries from the effects of the global financial cycle. In this sub-section, we examine whether an important policy variable—the stance of macroprudential regulation—has a similar effect. To this end, we estimate a version of column (7) in Table 2 where we separate the loadings to the PCs according to a country’s macroprudential stringency, and not its exchange rate regime. Specifically, we define a country as having tighter macroprudential regulations when it, cumulatively during our sample period, tightened its macroprudential policy stance more than the median country. As column (1) of Table 5 shows, the loadings to the second and third PC are substantially smaller and statistically insignificant for countries with tighter macroprudential regulation. For PC1, we find the loadings to be larger for countries with tighter macroprudential regualtions. As we show below, PC1 correlates strongly with the worldwide macroprudential strength, so that it is natural to see that PC1 loads more heavily on countries that follow the world trend more closely, i.e., have a tighter regulation. In columns (2)-(3), we split the overall sample into countries subject to a floating and fixed exchange rate, respectively, and obtain a similar picture, but the statistical significance declines with the smaller sample size.
Interestingly, we find that having a stricter macroprudential policy stance especially helps countries subject to fixed exchange rates insulate against PC2 and PC3, which together explain about 21% of variations in non-core ratios. For instance, the loading to PC2 (PC3) is 1.46 (1.56) for fixers with loose macroprudential regulation, whereas it is essentially zero for fixers with a tighter regulation. For countries on a floating exchange rate, macroprudential tightness does not provide any additional insulation device.

### v6

@@ 3 Empirical Approach
$u_{i,t}$ $=\boldsymbol{\omega}_{t}\left(\boldsymbol{\beta}_{i}\boldsymbol{\theta}_{i}^{x}+\boldsymbol{\theta}_{i}\right)+\boldsymbol{f}_{t}\left[\boldsymbol{\beta}_{i}\left(\boldsymbol{\gamma}_{i}^{x}+\boldsymbol{\theta}_{i}^{x}\boldsymbol{\xi}\right)+\boldsymbol{\gamma}_{i}+\boldsymbol{\theta}_{i}\boldsymbol{\xi}\right]+\epsilon_{i,t}$ (11)
and Equation (10) then becomes:
$y_{i,t}$ $=\rho_{i}y_{i,t-1}+\boldsymbol{\beta}_{i}\boldsymbol{\chi}_{i,t}+u_{i,t}.$ (12)
Our estimation approach consists of three steps. The first step is a CCE-MG estimation of (10) where the observed common factors are directly added and the cross-section averages span the remaining part of the factor space:
$y_{i,t}$ $=\rho_{i}y_{i,t-1}+\mathbf{x}_{i,t}\boldsymbol{\beta}_{i}+\boldsymbol{g}_{t}\boldsymbol{\theta}_{i}+\sum_{l=0}^{p}\bar{\boldsymbol{R}}_{t-l}\boldsymbol{\gamma}_{i,l}+e_{i,t}.$ (13)
An advantage of the CCE approach is that it incorporates the observed common factors in the same way as the observed country-specific variables. Therefore, as long as the number of cross-section averages exceeds the number of unobserved common factors, the coefficients $\boldsymbol{\beta}_{i}$ can be consistently estimated. The first step then approximates the unobserved common factors $\boldsymbol{f}_{t}$ with the cross-section averages, while the observed factors $\boldsymbol{g}_{t}$ are directly used. Implicitly, we are using the variation in $\mathbf{x}_{i,t}$, which is independent of the observed and unobserved common factors, $\boldsymbol{\chi}_{i,t}$, to obtain an estimate of $\boldsymbol{\beta}_{i}$. We then estimate the composite term $u_{i,t}$ from Equation (11):
$\hat{u}_{i,t}$ $=y_{i,t}-\hat{\rho}_{i}y_{i,t-1}-\mathbf{x}_{i,t}\boldsymbol{\hat{\beta}}_{i}.$ (14)
@@
Yet, our aim is the identification of the interaction of the observed common factors and the country-specific variation $\boldsymbol{\chi}_{i,t}$, like the exchange rate regime and other policy or regulatory parameters. To identify $\hat{u}_{i,t}$ correctly, we must ensure that no common factors or cross-section dependence is left in $\epsilon_{i,t}$.
@@
In the second step, we therefore extract the common components of the observed common factors, of the unobserved ones, and of the remaining residual variation from $\hat{u}_{i,t}$ by using principal components. The number of components is determined by the criteria of Ahn and Horenstein (2013). In the third step, we then replace the observed and unobserved common factors with the principal components from the previous step and estimate the following model:
@@
where $\boldsymbol{PC}_{t}$ are the first $p_{pc}$ principal components of the observed and unobserved common factors and of the estimated residuals and $\zeta_{i,t}$ is iid and uncorrelated across units. [^15] Our model is related to the regularized CCE estimator in Juodis (2022). The idea is to reduce the dimension of the cross-section averages to the dimension of the estimated number of common factors.
@@ 4.1 Benchmark Results
Real Interest Rate -0.039 -0.014 -0.020 -0.182 -0.313 -0.065 -0.065
@@
REER 0.098 0.106 0.067 0.704 0.659 0.424 0.424
(0.025)*** (0.025)*** (0.026)** (0.284)** (0.256)** (0.186)** (0.186)**
Capital-to-Asset -0.217 -0.216 -0.264 -0.135 -0.303 0.807 0.807
@@
This table displays the results for the years 2004-2022 with the non-core ratio as dependent variable. Standard errors in parenthesis. Significance levels: ${}^{*}10\%,^{**}5\%,^{***}1\%$. TWFE and TWFE-I include country and time fixed effects. TWFE-I additionally includes interactions between the fixed exchange rate dummy and time fixed effects. All explanatory variables except Capital Regulation and Liquidity Regulation are added as cross-section averages for MG-CCE. Principal Components (PC) are extracted from composite $\hat{u}_{i,t}$ in column MG-CCE. CD is the value of the CD test (Pesaran, 2015) with the null hypothesis of weak cross-section dependence. $\hat{M}_{X}$ is the estimate of the number of common factors using the Growth Criterion from Ahn and Horenstein (2013). $\hat{u}$ is the composite of observed and unobserved common factors from Equation (14) and $\hat{\epsilon}$ the residual. All variables are defined as in Table 9.
@@
Additional Results corresponding to columns (4), (5), (6) and (7) of Table 4.1. $F:PC=0$ tests if the coefficients of the principal components are joint significantly different from zero. $\Delta(all)$ test the null of slope homogeneity of all coefficients, $\Delta(PC)$ test the null of slope homogeneity of the coefficients of the principal components under the assumption that all other coefficients are heterogeneous. $PCF_{1}>PCF_{0}$ test if the sum of the coefficients of the PCs for $fixed=1$ are larger than those for $fixed=0$. Share PC is the share of the explained variance of the PC. All variables are defined as in Table 9.
@@
The first column of Table 4.1 presents a standard two-way fixed effects specification with homogeneous coefficients, country and year fixed effects, and no accounts of any observed or unobserved common factors. [^17] We use Stata 18 and the community contributed packages reghdfe (Correia, 2023) for the TWFE and FE estimations and xtdcce2 and xtcd2 (Ditzen, 2018, 2021) for the MG-CCE regressions and the CD tests. The number of common factors is estimated using xtnumfac (Ditzen and Reese, 2023). The CD test rejects the null of weak cross-section dependence, i.e., it suggests that there are global factors that are not accounted for and/or those global factors are affecting countries in a distinct way. [^18] Juodis and Reese (2021) show that the CD test statistic diverges under certain conditions and propose a corrected test. Our results are qualitatively similar using the their test. This is supported by two estimated common factors in the residuals ($\hat{M}_{\hat{\epsilon}}$). Moving rightwards along the table, we attempt to account for the factor structure in two ways. In column (2), we add the observed common factors and in column (3), we interact the time fixed effects with the fixed exchange rate indicator. Two of the observed common factors are now significant in column (2), but even in this case and in column (3) when interacting the time fixed effects with the exchange rate indicator, strong cross-section dependence remains in the residuals, indicating a potential bias of the regressions results.
@@
Moving to columns (4) and (5), we change our specification in two aspects. First, we allow for country-specific coefficients on both the country-specific variables and observed global factors using the mean group estimator (MG). The test for slope heterogeneity, presented in the columns labelled (4) and (5) in Table 4.1 confirms the necessity to allow for country-specific slopes. [^19] The tests for slope heterogeneity were performed with the xthst package (Bersvendsen and Ditzen, 2021). Turning to the results in column (4) of Table 4.1, while no factors remain in the residuals, the CD test still rejects the null of weak cross-section dependence. Further, none of the observed common factors are significant. While this specification potentially accounts for the factor structure, the results remain questionable due to the strong cross-section dependence. To overcome this weakness, in column (5), we then employ the CCE estimator as per equation (13)—stage (1) of our three step procedure—and add, in addition to the observed common factors, cross-section averages to fully account for the unobserved factor structure. [^20] We add the cross-section averages of the dependent variable, real exchange rate, GDP growth, the real interest rate, capital-to-asset ratio and money over GDP. We do not add the cross-section averages of the macroprudential variables as they are part of the observed common factor ’World Macroprudential’. The six cross-section averages ensure that the rank condition holds, i.e., the number of (un)observed factors is smaller or equal to the number of cross-section averages, see Karabiyik et al. (2017), DeVos et al. (2024). Together with the observed common factors, the cross-sectional averages span variables that are under direct policy control (like the exchange rate choice or macroprudential regulations) and those that are not under direct control (e.g., growth)., [^21] We also note that, in a very general setting, policy controlled variables in $x_{i,t}$ can interact with the common factors, resulting in higher order functions. In practice, these non-linear factors are linear in parameters and can be treated as additional factors in the estimation. An advantage of the CCE estimator is that we do not require knowledge of the exact factor structure. In the case of a time-invariant policy variable, such as the fixed or flexible exchange rate regime, the heterogeneous factor loadings soak up the interaction with the common factors. The same holds for interactions of the principal components and time-invariant policy variables when using a heterogeneous slope estimator. When using a pooled estimator, the interaction becomes part of the error term, resulting in strong cross-sectional dependence and biased estimates of the coefficients. The CD-test statistic falls into a non-rejection area (CD=0.797) and the estimate of the number of common factors in the residual is zero so that we can now calculate the composite $\hat{u}_{i,t}$, which encapsulates the entire factor structure. As we found three common factors in the non-core to core ratio, in the second step, we extract the first three principal components (PCs) from the composite $\hat{u}_{i,t}$. The PCs account for roughly 80% of the explained variance, see the lower panel in Table 4.1.222222We note that the estimate for the number of factors in the composite, $\hat{M}_{\hat{u}}$, is 2, the same as the number of factors in the residuals of the (TW)FE regressions. No factors remain in the residual, so we conclude that the composite captures the entire factor structure. Columns (6) and (7) of Table 4.1 present the results of the 3rd stage. Introducing the PCs successfully mitigates cross-sectional dependence, as indicated by the CD-statistic of -0.134 (p=0.894). As explained in the previous section, because the estimated PCs span both observed and unobserved global factors, their inclusion in the MG regression makes the observed common factors redundant (and they display high multicollinearity with some of them, as discussed below), so the latter are omitted. Hence we conclude that these two specifications fully account for the factor structure. In column (6), all three PCs are significant and positive, implying an increase in the non-core ratio following a shock propagated by the PCs. Column (7) then tests whether global factors (captured by the three PCs) affect the non-core ratio of countries on a fixed exchange rate more than those on a flexible exchange rate. The loadings for the first PC is higher in countries subject to fixed exchange rates (0.984 vs. 0.302) and significant. PC2 is positive and significant for fixers, while PC3 is significant for floaters. In general the loadings for fixers are larger than those for floaters, as confirmed by the test statistic of 7.12 on the differences between the sum of the PC loadings between fixed and flexible exchange rate regimes, shown in the column labelled (7) in Table 4.1. Therefore, the upshot is that there is cause for the trilemma hypothesis applied to banks’ non-core funding ratios.
@@
Across all specifications, it is mainly three domestic variables that affect banks’ non-core ratios in a statistically significant manner. First, the lagged dependent variable is positive and significant at the 1% level throughout. Second, banking sectors in countries with a more appreciated real exchange rate have higher non-core ratios. This result might reflect that more an appreciated real exchange rate makes it chaeper to borrow abroad. Finally, higher money-to-GDP ratios also tend to increase banks’ non-core ratios, consistent with the evidence on drivers of credit booms and the role of cross-border borrowing found in the studies referred to in previous sections.
@@ 4.2 Robustness Checks
In this sub-section, we show that our main results corresponding to column (7) of Table 4.1 are robust along several dimensions. We first drop the United States from our sample, as most of the observed common factors included in the regressions are based on (or at least dominated by) the US. Our benchmark results are hardly affected and we still find PC1 and PC3 to affect floaters, and PC1 and PC2 to affect fixers to a statistically significant extent, as column (1) of Table 4.2 shows.
In column (2), we use our benchmark sample including Iceland, which we dropped for the benchmark analysis because it is an extreme outlier for several of the variables included in the analysis. As becomes apparent, most results are similar to our baseline results, but the coefficients are estimated more imprecisely, at least for floaters, presumably due to the additional uncertainty introduced by the inclusion of Iceland. In addition, the specification is left with some cross-sectional dependence.
@@
Finally, in column (5), we use an alternative outcome variable—banks’ loan-to-deposit ratio. Previous work suggested it to be closely correlated with the non-core to core ratio, as we also confirm for our sample in Section 2. For this outcome variable, we find PC1 and PC2 to be significant drivers in countries with a flexible exchange rate and PC1 only to be statistically significant for countries on a fixed exchange rate. One difference to the previous results that we obtained for banks’ non-core ratios is that a flexible exchange rate does not provide any insulation to PC1. This result, however, is driven by the crises episodes included in column (5). In fact, when in unreported specifications we drop the Covid episode and/or the global financial crisis from the sample, we find that the loadings to PC1 are substantially larger for fixers than for floaters, as in our benchmark analysis.
@@
Real Interest Rate -0.093 -0.088 -0.094 0.066 -0.086
@@
This table displays the robustness checks. The non-core ratio is the dependent variable in Columns (1)-(4). In Column (5), LTD, the loan-to-deposit ratio, is the dependent variable. Standard errors in parenthesis. Significance levels: ${}^{*}10\%,^{**}5\%,^{***}1\%$. Column Excl. US excludes the United States. Column Inc. Iceland includes Iceland. Column No Covid spans the time period 2004:Q1-2019:Q4 and No Crisis 2010:Q1-2019:Q4. The 1st stage for each column is not displayed. The PCs are obtained from this initial stage. $\hat{M}_{X}$ is the estimate of the number of common factors using the Growth Criterion from Ahn and Horenstein (2013). $\hat{u}$ is the composite of observed and unobserved common factors from Equation (14) and $\hat{\epsilon}$ the residual. All variables are defined as in Table 9. For further details see notes of Table 4.1.
@@ 4.3 Does Macroprudential Policy Insulate Countries from the GFC?
Table 5: The Relevance of Macroprudential Policies
@@
Principal Components obtained from Baseline Regression in Table 4.1. The regressions in this table identify the loadings to the PCs separately for countries with loose vs. tight macroprudential policies. Column Fixed = 0 restricts the sample to countries with flexible exchange rates, Fixed = 1 to countries with fixed exchange rates. See the additional notes of Table 4.1. All variables are defined as in Table 9. For further details see notes of Tables 4.1 and 4.2.
Previous regressions have shown that a flexible exchange rate helps insulate countries from the effects of the global financial cycle. In this sub-section, we examine whether an important policy variable—the stance of macroprudential regulation—has a similar effect. To this end, we estimate a version of column (7) in Table 4.1 where we separate the loadings to the PCs according to a country’s macroprudential stringency, and not its exchange rate regime. Specifically, we define a country as having tighter macroprudential regulations when it, cumulatively during our sample period, tightened its macroprudential policy stance more than the median country. As column (1) of Table 4.2 shows, the loadings to the second and third PC are substantially smaller and statistically insignificant for countries with tighter macroprudential regulation. For PC1, we find the loadings to be larger for countries with tighter macroprudential regualtions. As we show below, PC1 correlates strongly with the worldwide macroprudential strength, so that it is natural to see that PC1 loads more heavily on countries that follow the world trend more closely, i.e., have a tighter regulation. In columns (2)-(3), we split the overall sample into countries subject to a flexible and fixed exchange rate, respectively, and obtain a similar picture, but the statistical significance declines with the smaller sample size.
Interestingly, we find that having a stricter macroprudential policy stance especially helps countries subject to fixed exchange rates insulate against PC2 and PC3, which together explain about 21% of variations in non-core ratios. For instance, the loading to PC2 (PC3) is 1.47 (1.55) for fixers with loose macroprudential regulation, whereas it is essentially zero for fixers with a tighter regulation. For countries on a flexible exchange rate, macroprudential tightness does not provide any additional insulation device.